OpenStudy (anonymous):

Calculate the distance between the points (1,5) and (2,-3). Round your answer to the nearest tenth. 8.1 65.0 8.3 31.4 1 points QUESTION 37 When calculating the midpoint of the points (3,1) and (11,7), what is the x-value of the midpoint? 3 7 11 14 1 points QUESTION 38 When calculating the midpoint of the points (3,1) and (9,7), what is the x-value of the midpoint? 1 3 6 7

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

sry, i'm busy. maybe @Directrix can help u

OpenStudy (anonymous):

its okay thank you

Directrix (directrix):

Try out the distance between two points calculator at this link. http://www.mathwarehouse.com/calculators/distance-formula-calculator.php The attachment shows you how the calculator looks. I'll also post the distance formula if you decide you want to use it.

Directrix (directrix):

Let me know what you get from the calculator if you chose to use it, okay? Attached is the distance formula if you prefer that.

OpenStudy (anonymous):

ok ill use it give me one min

Directrix (directrix):

It just acted the fool with me. I don't know what's up. See what happens with your computer.

OpenStudy (anonymous):

You can use simple distance/line/slope calculator: http://www.hackmath.net/en/calculator/line-slope?x0=1&y0=5&x1=2&y1=-3&submit=Solve Result is : The length of the segment AB: |AB| = 8.0623 Rounded to 10th is 8.1

Directrix (directrix):

@triangleguru has a better calculaor. Thanks.

Directrix (directrix):

To get the x-coordinate of the midpoint of the segment with endpoints (3,1) and (11,7), take the average of the two x-coordinates of the endpoints of the segment. So, that would be what?

OpenStudy (anonymous):

is it 7?

Directrix (directrix):

Let's see: (3+11)/2 = 14/2 = 7 YES, I agree.

Directrix (directrix):

On that third question, would you check to ensure that we are doing x-coordinate again? Somehow, I expected the question to ask what is the y-coordinate?

OpenStudy (anonymous):

were looking for the x-value

Directrix (directrix):

Okay, then do the same as on the previous question Take the average of the two x-coordinates.

OpenStudy (anonymous):

ok

Directrix (directrix):

Post what you get.

OpenStudy (anonymous):

i got 6

Directrix (directrix):

Let's see. (3+9)/2 = 12/2 = 6. YES, I agree. I have the midpoint formula attached in case you want to review it at some point.

OpenStudy (anonymous):

thank you for your help!

Directrix (directrix):

You are welcome.